OpenStudy (anonymous):

Medal will be awarded Patrick and Brooklyn are making decisions about their bank accounts. Patrick wants to deposit $300 as a principle amount, with an interest of 3% compounded quarterly. Brooklyn wants to deposit $300 as the principle amount, with an interest of 5% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

3% compounded quarterly \[300\times \left(1+\frac{.03}{4}\right)^{4\times 2}\]

OpenStudy (anonymous):

5% compounded monthly \[300\times \left(1+\frac{.05}{12}\right)^{2\times 12}\]

OpenStudy (anonymous):

i think the first method is better